If x=y-9, and y-4=10 x=?

Answer by ikleyn(52776)

.
x = y-9 = (y-4) - 5 = 10 - 5 = 5.

Answer: x = 5.


The lesson to learn from this solution:

        You do not need to solve the system to get the answer.